# On the Utility of Equivariance and Symmetry Breaking in Deep Learning Architectures on Point Clouds

### v1

@@ A.4 Additional Results and Discussion
Practical implications of the results Our experiments on ModelNet40 demonstrate similar trends to those discussed in the main paper for the experiments on ShapeNet, QM9 and CMU motion datasets. For a non-equivariant task (low geometric complexity) like classification, we see that equivariant methods perform best, model 6-7 in Tab . 4.
@@ A.5 Implementation details
ShapeNet 3D (Generation). For this task, we trained we trained model variation (1-4 & 16-19) in Tab.2 with two different settings of hidden features C = 256 (gray) and C = 2048. The later inflated model was trained to match the representation capacity of the rest of the models. For segmentation task, we use rotated samples and compute IOU with aligned and rotated samples. All the models were trained for 500 epochs with a learning rate $35e-$ 3 and weight decay of $81e-$ 8.
QM9 (Segmentation and Generation). For the segmentation task, we trained we trained model variation (1,2 & 6,7) in Tab.2 with two different settings of hidden features C = 256 (gray) and C = 2048. The later inflated model was trained to match the representation capacity of the rest of the models. All the models were trained for 500 epochs with a learning rate $35e-$ 3 and weight decay of $81e-$ 8.
CMU Motion Prediction. For this task we trained model variation (1-4 & 16-19) in Tab.3 with two different settings of hidden features C = 256 (gray) and C = 2048. The later inflated model was trained to match the representation capacity of the models (5-15). All the models were trained for 1000 epochs with a learning rate $35e-$ 3 and weight decay of $81e-$ 8.
@@
A.6 Additional figures
Illustrating the effect of equivariant models in ShapeNet segmentation
Figure 4: Depiction of an instance of ShapeNet segmentation for the class airplane. When the sample is rotated, a non-equivariant model mistakes wings for the nose of the airplane, while a rotationally equivariant model does the segmentation perfectly.
Human motion prediction using CMU motion capture dataset
Figure 5: Depiction of an instance from CMU motion capture dataset.

### v2

@@ A.5 Additional Results and Discussion
Practical implications of the results Our experiments on ModelNet40 demonstrate similar trends to those discussed in the main paper for the experiments on ShapeNet, QM9 and CMU motion datasets. For a non-equivariant task (low geometric complexity) like classification, we see that equivariant methods perform best, model 6-7 in Tab . 5.
@@ A.6 Implementation details
A.7 ShapeNet 3D (Segmentation and Generation).
ShapeNet 3D dataset (Chang et al., 2015) is used for part segmentation and generation tasks. ShapeNet consists of 16,881 shapes from 16 categories. Each shape is annotated with up to six parts, totaling 50 parts. We use the point sampling of 2,048 points and the train/validation/test split from (Qi et al., 2017).
For the generation task, we trained model variation (1-4 & 16-19) in Tab.4 and Tab.3 with two different settings of hidden features C = 256 (gray) and C = 2048. The inflated model was trained to match the representation capacity of the rest of the models. For segmentation task, we use rotated samples and compute IOU with aligned and rotated samples. All the models were trained for 500 epochs with a learning rate $35e-$ 3 and weight decay of $81e-$ 8.
A.8 QM9 (Generation).
QM9 dataset (Ramakrishnan R., 2014) contains up to 9 heavy atoms and 29 atoms including hydrogens. We use the train/val/test partitions introduced in Gilmer et al. (2017), which consists of 100K/18K/13K samples respectively, for each partition.
For the segmentation task, we trained model variation (1,2 & 6,7) in Tab.2 with two different settings of hidden features C = 256 (gray) and C = 2048. The inflated model was trained to match the representation capacity of the rest of the models. All the models were trained for 500 epochs with a learning rate $35e-$ 3 and weight decay of $81e-$ 8.
A.9 CMU Motion Prediction.
Figure 4: Depiction of an instance from CMU motion capture dataset.
For the motion prediction task, we evaluate our models on the CMU Human Motion Capture dataset (Gross & Shi, 2001), consisting of 31 equally connected nodes, each representing a specific position on the human body during walking. Given node positions at a random frame, the objective is to predict node positions after 30 time steps. As per Huang et al. (2022), we use the data of the 35th human subject for the experiment. See Figure 4 to see instance of the dataset. For this task we trained model variation (1-4 & 16-19) in Tab.4 with two different settings of hidden features C = 256 (gray) and C = 2048. The inflated model was trained to match the representation capacity of the models (5-15). All models were trained for 1000 epochs with a learning rate $35e-$ 3 and weight decay of $81e-$ 8.
@@
A.10 Additional figures
Rapidash architecture In the figure, we show a block design of Rapidash architecture with base space $\mathbb{R}^{3}$ (L) and $\mathbb{R}^{3}\times S^{2}$ (R).